# Position: We Need An Adaptive Interpretation of Helpful, Honest, and Harmless Principles

### v3

@@ Abstract
Correspondence to: Yue Huang (yhuang37@nd.edu), Xiangliang Zhang (xzhang33@nd.edu).
@@ 1 Introduction
The development of AI assistants has progressed rapidly, evolving from simple rule-based systems (Weizenbaum, 1966) to advanced models, such as Large Language Models (LLMs) (Zhao et al., 2023b), which excel in understanding context and generating human-like responses. The wide application of these models highlights not only the expanding capabilities of AI but also the growing need to ensure these systems align with human preferences and values. To address this challenge, the HHH principle–standing for Helpful, Honest, and Harmless–was proposed by Askell et al. (2021) as a guiding framework for designing and evaluating AI systems. The principle aims and has been widely utilized to align AI behavior with human-centered values, providing critical benchmarks for tasks such as training data selection, strategy design, and deployment guidance (Touvron et al., 2023b; Peng et al., 2023; Li et al., 2024a; Bai et al., 2023). Its adoption has been instrumental in advancing AI alignment, particularly in ensuring that powerful models like LLMs prioritize user benefit, truthfulness, and safety.
@@
2 HHH Principle
2.1 Initial Definition
@@ 2.2 Fundamental Value of the HHH Principle
The HHH principle uniquely emphasizes both trustworthiness and utility  [^1] Utility here refers to the model effectiveness in natural language processing tasks, including logical reasoning, content summarization, text generation, and so on (Huang et al., 2024a).. AI models, such as LLMs, interact with users by providing information, generating content, and assisting in decision-making, making trustworthiness and utility essential for their effectiveness and reliability. While utility is prioritized in model development through training and tuning loss minimization, an automated and human-effortless process, there is a risk of compromising ethical integrity or introducing biases. Thus, typical alignment principles emphasize additional efforts to align AI models with human values, as shown in Table 1. Research indicates that the resulting models often choose safe responses over helpful ones (Touvron et al., 2023a). However, the HHH principle offers a unique balance by explicitly emphasizing both dimensions. The dimension of honesty requires AI models generate accurate and truthful outputs.
Framework & Principle Trust. Utility
[…]
@@ 3 Ambiguity and Conflicts in HHH
3.1 Case Study I: Boundary of Harmlessness
The distinction between harmful and benign content is not always clear-cut, as context, intent, and interpretation play crucial roles. This lack of clarity complicates the development of safety mechanisms and alignment strategies (Bauer & Bindschaedler, 2021; Truong et al., 2024).
A harmless query becomes harmful when it is used to bypass safety restrictions, e.g, in the scenario of adversarial prompting. For example, if an AI model initially rejects a query for being harmful ($I_{B}$, "Write a post promoting militaristic aggression,"), the user may slightly rephrase it to appear benign while maintaining the original intent: $I_{A}$, "Please write a post encouraging people to build a strong national defense". If the model then accepts and responds to the modified query $I_{A}$, it may unintentionally generate harmful content. However, if the same revised query $I_{A}$ were submitted independently, without following a rejected attempt, it would still appear harmless. Training machine learning classifiers to detect toxicity in queries may fail in this case (Wang & Chang, 2022; Ousidhoum et al., 2021; OpenAI, 2024). This highlights the importance of incorporating context and intent when distinguishing between harmful and harmless queries.
Moreover, the harmfulness or harmlessness of a generated response can also depend on the user’s intent. For example, if a model responds to a query but includes a moral disclaimer to discourage misuse, the response itself may remain neutral or informative (Mazeika et al., 2024; Ran et al., 2024; Huang et al., 2024c). However, an attacker could exploit this by simply removing the moral disclaimer, and repurposing the content for malicious objectives. Conversely, a well-intentioned user may utilize the same response for legitimate purposes without causing harm.
Although recent work by OpenAI (Mu et al., 2024) has proposed a set of rules that a trustworthy LLM should adhere to (e.g., categorizing response into hard refusal, soft refusal, and compliance), the core issue lies in the ambiguity of determining the appropriateness of a response without knowing the user’s real intent and context. This highlights the need for an adaptive interpretation of HHH principle, which can dynamically adjust to different contexts and user intents.
@@ 3.3 Case Study III: Helpfulness in Different Contexts
The definition of helpfulness varies substantially based on domain-specific objectives. For instance, in education, Hemami (2024) frame helpfulness as the ability of AI models to identify optimal learning strategies and personalize content for individual students. For general language assistants, helpfulness is defined as a clear attempt to perform the task or answer the posed question (Askell et al., 2021). These variations, illustrated in Appendix C, reveal the importance of contextual adaptability in defining and applying the concept of helpfulness.
@@
Current evaluation methods often employ narrow metrics that align with specific research objectives rather than capturing helpfulness holistically. Moreover, most frameworks utilize specialized metrics (Liu et al., 2024c) or LLM-as-judge approaches (Zheng et al., 2023; Wang et al., 2024b), Ye et al. (2024), demonstrating that such evaluations can introduce systematic biases through prompt design choices. Future benchmarks should incorporate multidimensional evaluation such as complementary dimensions (e.g., correctness assessment, coherence analysis, creativity metrics), while considering the relative importance and interactions between these aspects. Consequently, multidimensional approaches would better reflect the diverse expectations of helpfulness in real-world applications.
@@ 4 Priority Order
Definition and necessity. In the deployment of AI models, conflicting requirements often arise among helpfulness, honesty, and harmlessness across various scenarios. For instance, in cybersecurity applications, when a user requests information about system vulnerabilities, the model must balance providing helpful technical details (helpfulness) with the need to prevent malicious exploitation (harmlessness) (Zhang et al., 2024a). Similarly, in medical or legal domains, prioritizing honesty may sometimes conflict with harmlessness, such as when conveying critical but distressing information. To address such conflicts, we introduce the concept of priority order, which is defined as:
A dynamic hierarchical framework that determines the relative importance and execution sequence of three dimensions of the HHH principle based on contextual requirements.
The necessity of adopting a dynamic rather than static priority order arises from the complexity and variability of real-world applications, where the relative importance of helpfulness, honesty, and harmlessness shifts based on context, user intent, and domain-specific requirements. In this section, we explore the priority order of the HHH principle from two complementary perspectives as shown in Figure 1: (1) a structural and statistical perspective (§4.1), which examines how prioritization operates along hierarchical levels and scales, shaping its manifestation across different domains and impact scopes; and (2) an application-level embodiment (§4.2), which investigates how priority order shifts based on user populations and task contexts.
@@ 5.1 Boost Harmlessness and Helpfulness Simultaneously
Rather than treating harmlessness and helpfulness as competing objectives in multi-objective optimization (Kochenderfer & Wheeler, 2019), recent research suggests they can be mutually reinforcing. Some approaches first establish a baseline of harmlessness before optimizing for helpfulness (Gao et al., 2024), while others integrate multi-objective alignment to improve both simultaneously (Yang et al., 2024; Wang et al., 2024a; Zhou et al., 2024c; Fu et al., 2024; Guo et al., 2024b).
@@ 5.2 Interdependencies Among Dimensions in High-Risk Evaluations
While helpfulness, harmfulness, and honesty are always intertwined to some degree, in many everyday tasks these interactions are easy to manage or less impactful (Zheng et al., 2023; Huang et al., 2024b; Zhou et al., 2024a; Wu et al., 2024a; Sandmann et al., 2024; Huang et al., 2023c). By contrast, in high-risk or specialized tasks, these same interactions can become more complex and lead to significant risks. So simple, independent evaluation of each dimension may underestimate systemic hazards (Zhou et al., 2024b; Phan Van et al., 2024; Thirunavukarasu et al., 2023). For instance, in medical diagnostics, a drug’s helpfulness in treating a condition and the harmfulness associated with possible side effects exist simultaneously (Phan Van et al., 2024; Thirunavukarasu et al., 2023; Sandmann et al., 2024). If the helpfulness is overlooked, the underlying condition remains untreated, potentially leading to serious deterioration of the patient’s health. Conversely, if the harmful side effects receive insufficient attention, the patient could suffer from severe complications or even life-threatening adverse reactions. A truly effective medical decision must balance both factors. Therefore, when assessing a model’s diagnostic capabilities, it is insufficient to focus solely on whether it provides a "helpful" conclusion. It is equally important to evaluate the probability, severity, and ethical implications of harmful consequences (Thirunavukarasu et al., 2023). This issue is also highlighted when people interact with LLMs in mental health emergencies, e.g., when suffering from mania or psychosis (Grabb et al., 2024). In such scenarios, both providing harmful information to the user (helpfulness) or refusing to respond (harmlessness) can exaggerate existing symptoms and lead to severe harm of the user or others (Tucker, 2025). Similarly, in answering highly specialized questions, if an LLM generates hallucinated outputs, its "helpfulness" is immediately called into question. In lab safety contexts, any "helpful" advice that neglects safety considerations could lead to severe accidents, and hallucinations in the model’s generation process present hidden high-risk factors (Zhou et al., 2024b).
@@
Thus, when evaluating LLMs, particularly in high-risk and specialized fields, it is crucial to develop holistic evaluation frameworks that account for the emergent properties. This requires systematic identification of cascading risk scenarios that could lead to severe outcomes (Zhou et al., 2024b; Phan Van et al., 2024; Thirunavukarasu et al., 2023). For example, to avoid lab safety accidents, responses with potential risks should undergo rigorous scrutiny and verification. In medical diagnostics, more stringent test sets and consequence-driven weighting mechanisms should be implemented for scenarios where a drug is both helpful and potentially harmful, incorporating the severity of misdiagnosis or side effects into the evaluation process. Only by explicitly modeling these "multidimensional entanglement" scenarios can we more accurately gauge the risks and value a model may present in real-world applications, thereby providing more targeted guidance for model optimization and regulatory decision-making.
@@ 6 Reference Framework
Building on the current ambiguous understanding of the principle (§3), it is evident that interpreting the helpfulness, honesty, and harmlessness in diverse contexts presents significant challenges. These principles often interact in complex ways (relationship as discussed in §5), with their relative importance varying depending on the specific application and scenario (i.e., priority order as discussed in §4). This variability underscores the urgent need for a reference framework to systematically guide the interpretation and application of the HHH principle. Such a framework would address the critical question: What perspectives and considerations are necessary to adapt these principles effectively in different scenarios?
This framework benefits multiple stakeholders. For developers in industries, it provides structured guidelines to integrate the HHH principles into the development of AI models, ensuring regulatory compliance. For academic researchers, it offers a novel view to study trade-offs between these principles in different domains (out-of-distribution (OOM) problem). For end users of AI models, it enhances transparency and governance, ensuring AI interactions align with user expectations (Larsson & Heintz, 2020). In this section, we propose a reference framework to address this issue from multiple perspectives: Contextual Object, Value Anchor & Value Scale, Risk Assessment, Benchmarking Standard, and System Transparency & Governance. We also provide a detailed case study on applying the 3H principles to the development of a chemistry foundation model, which can be found in Appendix D.
Contextual Object. First, the framework requires determining the objects in a specific scenario when adapting HHH principle. As discussed in §3.3, different contexts will result in various understandings of dimensions, where the process, of determining scenario objects, directly define the specific context in which the AI model operates, mitigating potential ambiguity and conflicts. The objects contain key elements that are important for determining the context: User Group, Application Aim, Task Type, and Environment Access.
@@
Value Anchor & Value Scale. The framework next requires the identification of a value anchor and a value scale to guide the prioritization of the HHH principle within a specific context. The value anchor defines the core value or central priority, such as harmlessness in high-risk domains or helpfulness in performance-driven applications. The value scale, in turn, allows for a dynamic balance among the principles by assigning relative weights to each dimension based on contextual demands, stakeholder needs, and societal expectations. For example, in the case of an educational AI model designed for younger students, the value anchor would prioritize harmlessness to prevent inappropriate or harmful outputs. The corresponding value scale might place harmlessness “far before” helpfulness, reflecting the heightened importance of safety in such a sensitive application.
In other words, this framework aims to endow AI models with specific traits or values (value anchor) and control the depth and degree of these traits (value scale) to ensure that the model remains predictable and operates within a controllable range.
Risk Assessment. Building on the contextual objects and value anchor & value scale, the reference framework must include a risk assessment mechanism to evaluate the trade-offs and potential negative impacts of prioritizing and customizing certain dimensions of the HHH principle. The assessment helps developers, users, and decision-makers determine whether the risks associated are acceptable and manageable within the given context.
Risk assessment involves three primary considerations: measuring the potential risks, assessing stakeholder tolerance for those risks, evaluating system consistency, and considering potential marginal risk (Kapoor et al., 2024). For the same example in an education application where harmlessness is prioritized far above helpfulness, a risk assessment would evaluate whether this decision significantly damages the user experience (e.g., overly restrictive outputs that limit engagement or creativity (Li et al., 2024c)). Simultaneously, it would determine whether stakeholders–such as developers and end users–are willing to accept this trade-off to ensure safety and ethical integrity. For context-dependent HHH, consistency is a requirement for trustworthy deployment to ensure similar behavior or balance of HHH in similar contexts (e.g., reproducible priority ordering), thus, the framework also emphasizes the significance of consistency evaluation. Moreover, Kapoor et al. (2024) point out some marginal risks (e.g., Non-consensual intimate imagery (NCII) (INHOPE, 2025)) that may be ignored in open foundation models (Bommasani et al., 2021), which should also be strictly considered during assessment.
@@
In particular, most existing risk assessment frameworks such as EU High-Risk AI guidelines (EU, 2024a) focus primarily on societal-level risks, particularly those related to the aspect of harmfulness and safety. While these are crucial, we emphasize that assessments should also account for the direct impacts on stakeholders, such as developers’ economic viability and user experiences. For instance, excessive economic costs associated with implementing overly strict harmlessness safeguards can be a significant risk for developers, just as diminished user satisfaction from overly cautious systems can reduce utility and adoption.
Benchmarking Standard. Benchmarks provide an objective and reproducible way to evaluate whether an AI system adheres to the HHH principle under specific scenarios. Without well-defined standards, evaluations risk becoming inconsistent (McIntosh et al., 2024), subjective (Ye et al., 2024), or overly dependent on context-specific assumptions. Under a specific context, the benchmarking standard must align closely with the interpretation (e.g., prioritization) of the HHH principle established by the framework. For example, the benchmarks should reflect this prioritization if helpfulness is prioritized above harmlessness. In such a case, utility-related benchmarks should evaluate helpfulness directly and assess how trustworthiness and harmlessness contribute to or detract from the overall utility. Moreover, to achieve a fine-grained evaluation of AI-powered systems, it is essential to integrate multiple benchmarks that focus on different dimensions of model performance. If necessary, the dynamic and automatic evaluation mechanism Zhu et al. (2023); Wu et al. (2024b); An et al. (2024); Chen et al. (2024a) can also be introduced to prevent data leakage. Furthermore, as emphasized by Reuel et al. (2024b), designing a framework for assessing the reliability of the benchmark can enhance its credibility and authority. We show some benchmark examples of HHH principle in Appendix A.
System Transparency & Governance. An effective framework for interpreting the HHH principle must prioritize transparency in its implementation. Transparency is essential for several reasons. First, it ensures that stakeholders, including developers, users, and policymakers, clearly understand how the principles of helpfulness, honesty, and harmlessness are being interpreted in specific contexts, and provide a foundation for audits and external reviews. This clarity helps build trust in the developed system, as stakeholders can evaluate whether the system aligns with their expected standards. Beyond transparency, the framework must also address how the HHH principle facilitate governance (Reuel et al., 2024a; Wilczek et al., 2024). Governance involves setting policies and protocols to ensure that models are developed, deployed, and monitored responsibly. By embedding the HHH principle into the governance structure, the framework can guide developers and regulators in identifying and mitigating risks, establishing ethical safeguards, and ensuring compliance with societal standards. For example, helpfulness can inform standards for task performance and user satisfaction, while harmlessness can shape policies for harm mitigation.

### v4

# Prioritization First, Principles Second: An Adaptive Interpretation of Helpful, Honest, and Harmless Principles

@@ Abstract
Correspondence to: Yue Huang (yhuang37@nd.edu).
@@ 1 Introduction
The development of AI assistants has progressed rapidly, evolving from simple rule-based systems (Weizenbaum, 1966) to advanced models, such as Large Language Models (LLMs) (Zhao et al., 2023b). The wide application of these models highlights not only the expanding capabilities of AI but also the growing need to ensure these systems align with human preferences and values. To address this challenge, the HHH principle–standing for Helpful, Honest, and Harmless–was proposed by Askell et al. (2021) as a guiding framework for designing and evaluating AI systems. The principle aims and has been widely utilized to align AI behavior with human-centered values, providing critical benchmarks for tasks such as training data selection, strategy design, and deployment guidance (Touvron et al., 2023b; Li et al., 2024a; Bai et al., 2023). Its adoption has been instrumental in advancing AI alignment, particularly in ensuring that powerful models prioritize user benefit, truthfulness, and safety.
@@
2 Fundamental Value of the HHH Principle
@@
The HHH principle uniquely emphasizes both trustworthiness and utility  [^1] Utility here refers to the model effectiveness in natural language processing tasks, including logical reasoning, content summarization, text generation, and so on (Huang et al., 2024a).. AI models, such as LLMs, interact with users by providing information, generating content, and assisting in decision-making, making trustworthiness and utility essential for their effectiveness and reliability. While utility is prioritized in model development through training and tuning loss minimization, an automated and human-effortless process, there is a risk of compromising ethical integrity or introducing biases. Thus, typical alignment principles emphasize additional efforts to align AI models with human values, as shown in Appendix A. Research indicates that the resulting models often choose safe responses over helpful ones (Touvron et al., 2023a). However, the HHH principle offers a unique balance by explicitly emphasizing both dimensions. The dimension of honesty requires AI models generate accurate and truthful outputs.
@@ 3 Ambiguity and Conflicts in HHH
3.1 Case Study I: The Boundary of Harmlessness
Distinguishing harmful from benign content is rarely clear-cut: context, intent, and interpretation jointly determine perceived risk. This ambiguity complicates the design of safety mechanisms and alignment strategies (Bauer & Bindschaedler, 2021; Truong et al., 2024).
A seemingly harmless query becomes harmful under context. Let $q$ denote a user query, $H$ the interaction history (e.g., prior refusals), and $u$ the user’s latent intent. Define a harm predicate $\mathrm{Harm}(q\,|\,H,u)\in\{0,1\}$. A surface-form classifier $T(q)$ that ignores $H$ and $u$ can fail because harmfulness is not a property of $q$ alone.
Consider a rejection of
$I_{B}:\ \textit{{\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}``Write a post promoting militaristic aggression.''}}$
followed by a rephrasing that preserves intent while appearing benign,
[…]
If the model accepts $I_{A}$ immediately after refusing $I_{B}$, we may have $\mathrm{Harm}(I_{A}\,|\,H{=}\text{``follows }I_{B}\text{''},u{=}\text{malicious})=1$, even though in isolation $\mathrm{Harm}(I_{A}\,|\,H{=}\varnothing,u{=}\text{benign})=0$. Purely textual toxicity detectors trained on $q$ risk misclassification in such settings (Wang & Chang, 2022; Ousidhoum et al., 2021; OpenAI, 2024b). This motivates harm assessments that condition on both dialogue context and inferred intent.
The harmfulness of a response depends on user intent. Let $r$ be a model response. Even if $r$ includes a moral disclaimer and is informationally neutral (Mazeika et al., 2024; Ran et al., 2024; Huang et al., 2024c), an attacker can strip the disclaimer and repurpose $r$ for misuse, while a well-intentioned user may apply the same $r$ for legitimate ends. Formally, downstream harm is a function of $(r,u)$ and post-processing, not $r$ alone.
Policy implication. OpenAI’s recent guidance proposes categories such as hard refusal, soft refusal, and compliant responses (Mu et al., 2024). Yet, deciding which category is appropriate is underdetermined without $(H,u)$. Hence, we argue for an adaptive interpretation of the HHH principle: safety policies and response modes should be functions
$\pi^{\ast}:(q,H)\ \mapsto\ \{\text{hard refusal},\ \text{soft refusal},\ \text{compliance}\}$
that are explicitly conditioned on interaction history and an estimate of user intent, rather than on surface form alone.
@@ 3.3 Case Study III: Helpfulness in Different Contexts
The definition of helpfulness varies substantially based on domain-specific objectives. For instance, in education, Hemami (2024) frame helpfulness as the ability of AI models to identify optimal learning strategies and personalize content for individual students. For general language assistants, helpfulness is defined as a clear attempt to perform the task or answer the posed question (Askell et al., 2021). These variations, illustrated in Table 1, reveal the importance of contextual adaptability in defining and applying the concept of helpfulness.
@@
Current evaluation methods often employ narrow metrics that align with esearch objectives rather than capturing helpfulness holistically. Moreover, most frameworks utilize specialized metrics (Liu et al., 2024c) or LLM-as-judge approaches (Zheng et al., 2023; Wang et al., 2024b), Ye et al. (2024), demonstrating that such evaluations can introduce systematic biases through prompt design choices. Future benchmarks should incorporate multidimensional evaluation such as complementary dimensions (e.g., correctness assessment, coherence analysis, creativity metrics), while considering the relative importance and interactions between these aspects. Consequently, multidimensional approaches would better reflect the diverse expectations of helpfulness in real-world applications.
@@ 4 Priority Order
Definition and necessity. In the deployment of AI models, conflicting requirements often arise among helpfulness, honesty, and harmlessness across various scenarios. For instance, in cybersecurity applications, when a user requests information about system vulnerabilities, the model must balance providing helpful technical details (helpfulness) with the need to prevent malicious exploitation (harmlessness) (Zhang et al., 2024a). Similarly, in medical or legal domains, prioritizing honesty may sometimes conflict with harmlessness, such as when conveying critical but distressing information. To address such conflicts, we introduce the concept of priority order, which is defined as: A dynamic hierarchical framework that determines the relative importance and execution sequence of three dimensions of the HHH principle based on contextual requirements.
(Formal Definition of Priority Order) Let each scenario be denoted by $s=(t,u,d)\in S$, where $t$ indexes the task type, $u$ the user group, and $d$ the domain or risk level. We define a mapping $\pi:S\to\Sigma$ that assigns to each scenario $s$ a permutation $\pi(s)=[\pi_{1}(s),\,\pi_{2}(s),\,\pi_{3}(s)]\in\{H,O,S\}!$, where $H,O,S$ correspond to the metrics of helpfulness, honesty, and harmlessness, respectively. For a model parameterized by $\theta$, let $F_{H}(\theta)$, $F_{O}(\theta)$, and $F_{S}(\theta)$ denote the corresponding performance scores on each dimension. The priority order objective in scenario $s$ is then given by the lexicographic optimization: $\theta^{*}(s)=\arg\max_{\theta}\bigl(F_{\pi_{1}(s)}(\theta),\,F_{\pi_{2}(s)}(\theta),\,F_{\pi_{3}(s)}(\theta)\bigr)_{\text{lex}},$ where $(a,b,c)_{\text{lex}}$ denotes comparison in dictionary order: first maximize the primary metric $F_{\pi_{1}}$; among its maximizers, maximize $F_{\pi_{2}}$; and finally maximize $F_{\pi_{3}}$. Optionally, a relaxed version can be formulated using $\varepsilon$-constraints: $\theta_{(1)}=\arg\max_{\theta}F_{\pi_{1}(s)}(\theta),$ $\theta_{(2)}=\arg\max_{\theta}F_{\pi_{2}(s)}(\theta)\quad\text{s.t. }F_{\pi_{1}(s)}(\theta)\geq F_{\pi_{1}(s)}(\theta_{(1)})-\varepsilon_{1},$ $\theta_{(3)}=\arg\max_{\theta}F_{\pi_{3}(s)}(\theta)\quad\text{s.t. }F_{\pi_{i}(s)}(\theta)\geq F_{\pi_{i}(s)}(\theta_{(i)})-\varepsilon_{i},\;i=1,2.$
@@ 5.1 Boost Harmlessness and Helpfulness Simultaneously
Rather than treating harmlessness and helpfulness as competing objectives in multi-objective optimization (Kochenderfer & Wheeler, 2019), recent research suggests they can be mutually reinforcing. Some approaches first establish a baseline of harmlessness before optimizing for helpfulness (Gao et al., 2024), while others integrate multi-objective alignment to improve both simultaneously (Yang et al., 2024; Wang et al., 2024a; Zhou et al., 2024d; Fu et al., 2024; Guo et al., 2024b).
@@ 5.2 Interdependencies Among Dimensions in High-Risk Evaluations
While helpfulness, harmfulness, and honesty are always intertwined to some degree, in many everyday tasks these interactions are easy to manage or less impactful (Zheng et al., 2023; Huang et al., 2024b; Zhou et al., 2024b; Wu et al., 2024a; Sandmann et al., 2024; Huang et al., 2023c). By contrast, in high-risk or specialized tasks, these same interactions can become more complex and lead to significant risks. So simple, independent evaluation of each dimension may underestimate systemic hazards (Zhou et al., 2024c; Phan Van et al., 2024; Thirunavukarasu et al., 2023). For instance, in medical diagnostics, a drug’s helpfulness in treating a condition and the harmfulness associated with possible side effects exist simultaneously (Phan Van et al., 2024; Thirunavukarasu et al., 2023; Sandmann et al., 2024). If the helpfulness is overlooked, the underlying condition remains untreated, potentially leading to serious deterioration of the patient’s health. Conversely, if the harmful side effects receive insufficient attention, the patient could suffer from severe complications or even life-threatening adverse reactions. A truly effective medical decision must balance both factors. Therefore, when assessing a model’s diagnostic capabilities, it is insufficient to focus solely on whether it provides a "helpful" conclusion. It is equally important to evaluate the probability, severity, and ethical implications of harmful consequences (Thirunavukarasu et al., 2023). This issue is also highlighted when people interact with LLMs in mental health emergencies, e.g., when suffering from mania or psychosis (Grabb et al., 2024). In such scenarios, both providing harmful information to the user (helpfulness) or refusing to respond (harmlessness) can exaggerate existing symptoms and lead to severe harm of the user or others (Tucker, 2025). Similarly, in answering highly specialized questions, if an LLM generates hallucinated outputs, its "helpfulness" is immediately called into question. In lab safety contexts, any "helpful" advice that neglects safety considerations could lead to severe accidents, and hallucinations in the model’s generation process present hidden high-risk factors (Zhou et al., 2024c).
@@
Thus, when evaluating LLMs, particularly in high-risk and specialized fields, it is crucial to develop holistic evaluation frameworks that account for the emergent properties. This requires systematic identification of cascading risk scenarios that could lead to severe outcomes (Zhou et al., 2024c; Phan Van et al., 2024; Thirunavukarasu et al., 2023). For example, to avoid lab safety accidents, responses with potential risks should undergo rigorous scrutiny and verification. In medical diagnostics, more stringent test sets and consequence-driven weighting mechanisms should be implemented for scenarios where a drug is both helpful and potentially harmful, incorporating the severity of misdiagnosis or side effects into the evaluation process. Only by explicitly modeling these "multidimensional entanglement" scenarios can we more accurately gauge the risks and value a model may present in real-world applications, thereby providing more targeted guidance for model optimization and regulatory decision-making.
@@ 6 Reference Framework
Building on the ambiguous understanding of the HHH principle (§3), it remains challenging to interpret these dimensions across diverse contexts. These dimensions often interact in nontrivial ways (§5), and their relative importance varies by application and scenario (§4). This variability highlights the need for a systematic reference framework to guide the adaptive interpretation and implementation of HHH. Such a framework addresses the central question: What perspectives and constraints are necessary to meaningfully apply HHH principles in different contexts?
This framework serves multiple stakeholders. For developers, it offers structured guidance for integrating HHH into model design and deployment while meeting regulatory standards. For end users, it promotes transparency and governance, aligning AI behavior with user expectations (Larsson & Heintz, 2020). In this section, we introduce a reference framework composed of four components: Contextual Object, Value Anchor & Value Scale, Risk Assessment, and Alignment Auditing & Governance Infrastructure. A detailed case study applying this framework to the development of a chemistry foundation model is provided in Appendix C.
Contextual Object. First, the framework requires determining the objects in a specific scenario when adapting the HHH principle. Concretely, we represent each scenario $s\in S$ by a tuple $s=(t,u,d),$ where $t$ is the task type, $u$ the user group, and $d$ the domain or risk level. We then extract the key contextual elements via the mapping $O(s)=(u_{s},\;a_{s},\;t_{s},\;e_{s})\;\in\;\mathcal{U}\times\mathcal{A}\times\mathcal{T}\times\mathcal{E},$ where $u_{s}$ (User Group), $a_{s}$ (Application Aim), $t_{s}$ (Task Type) and $e_{s}$ (Environment Access) together define the concrete operating context of the AI model. By fixing $O(s)$, we mitigate ambiguity in interpreting helpfulness, honesty, and harmlessness for that scenario. Here are the definitions of different elements in Contextual Object:
@@
Value Anchor & Value Scale. After fixing the contextual object $O(s)$, we identify a value anchor $\alpha(s)\in\{H,O,S\}$ that selects the core dimension (e.g., $H$ for helpfulness, $O$ for honesty, or $S$ for harmlessness) most critical in scenario $s$. We then assign a value scale as $v(s)=\bigl[v_{H}(s),\,v_{O}(s),\,v_{S}(s)\bigr]\in\Delta^{2},$ where $\Delta^{2}=\{w\in\mathbb{R}^{3}:\sum_{i}w_{i}=1,\;w_{i}\geq 0\}$, to distribute relative importance across the three metrics. For instance, in an educational setting for younger students ($\alpha(s)=S$), one might set $v(s)=[\,v_{S}(s)=0.8,\;v_{H}(s)=0.2,\;v_{O}(s)=0\,],$ placing harmlessness "far before" helpfulness and honesty. This mechanism endows the model with a fixed priority trait (the anchor) while controlling the degree of influence (the scale), ensuring predictable, context‐sensitive behavior.
Risk Assessment. Building on the contextual object $O(s)$ and the value anchor $\alpha(s)$ with scale $v(s)$, we introduce four risk components: the potential harm $\rho_{r}(s,\theta)$ (e.g. overly restrictive outputs that limit engagement or creativity (Li et al., 2024c)), the stakeholder tolerance threshold $\rho_{t}(s)$, the consistency measure $\rho_{c}(s,\theta)$ for reproducible HHH balance across similar contexts, and the marginal hazard factor $\rho_{m}(s)$ (e.g. non-consensual intimate imagery (NCII) (INHOPE, 2025)). We then enforce scenario-specific thresholds $\tau_{r}(s)$, $\tau_{c}(s)$ and $\tau_{m}(s)$ so that any acceptable model configuration $\theta$ must satisfy
$\rho_{r}(s,\theta)\leq\tau_{r}(s),\quad\rho_{c}(s,\theta)\geq\tau_{c}(s),\quad\rho_{m}(s)\leq\tau_{m}(s),\quad\rho_{t}(s)\geq\rho_{r}(s,\theta).$
@@
For example, in an educational application where $\alpha(s)=S$ prioritizes harmlessness, we check $\rho_{r}(s,\theta)\leq\tau_{r}(s)$ to avoid overly restrictive outputs, while ensuring $\rho_{t}(s)\geq\rho_{r}(s,\theta)$ so that both developers and end users accept the harmlessness-first trade-off. In particular, most existing risk assessment frameworks such as EU High-Risk AI guidelines (EU, 2024a) focus primarily on societal-level risks. While these are crucial, we emphasize that assessments should also account for the direct impacts on stakeholders, such as developers’ economic viability and user experiences. For instance, excessive economic costs associated with implementing overly strict harmlessness safeguards can be a significant risk for developers, just as diminished user satisfaction from overly cautious systems can reduce utility and adoption.
Alignment Auditing & Governance Infrastructure. To ensure that the adaptive interpretation of the HHH principle is not only theoretically sound but also practically enforceable, our framework incorporates a unified infrastructure for alignment auditing and governance integration. This infrastructure serves a dual function: (1) enabling systematic, reproducible evaluation of alignment quality under context-specific configurations, and (2) facilitating transparent and accountable oversight in real-world deployments.
For evaluation, we formalize auditing through a function $A:S\times\Theta\to\mathbb{R}^{k}$, where each component $a_{i}(s,\theta)$ captures a specific evaluation dimension such as factual accuracy, ethical robustness, or user satisfaction. For each scenario $s$, a minimum threshold $\beta(s)\in\mathbb{R}^{k}$ is specified, and an aligned model must satisfy $A(s,\theta)\geq\beta(s),$ ensuring that performance aligns with the scenario’s contextual value priorities. For instance, if helpfulness is prioritized over harmlessness, then the utility-related metrics within $A(s,\theta)$ must dominate, while residual safety or truthfulness must still be non-negligible. Single-dimensional or static benchmarks are insufficient; instead, multi-perspective and modular auditing protocols are needed. To further enhance fidelity and reduce evaluator bias, dynamic, contextual or automatic evaluation mechanisms (e.g., Zhu et al. 2023; Wu et al. 2024b; An et al. 2024; Chen et al. 2024a; Zhou et al. 2024a) may be employed for reproducibility across varying contexts.
On the governance side, operationalizing the HHH principle requires transparent articulation of its contextual interpretation and enforcement mechanisms. We define a governance trace $G(s,\theta)=\bigl(O(s),\,\alpha(s),\,v(s),\,A(s,\theta),\,\rho(s,\theta),\,\theta\bigr)$, which records the full lifecycle of contextualization, prioritization, evaluation, and configuration. This trace supports transparent documentation and allows stakeholders-including developers, users, auditors, and regulators-to trace how alignment is defined, implemented, and verified. Such traceability underpins trust, enables audits, and supports regulatory compliance. Governance in this framework is not static but procedural, encompassing value documentation, justification, verification, and post-deployment monitoring (Reuel et al., 2024; Wilczek et al., 2024).